# Optimal Transport for $\epsilon$-Contaminated Credal Sets

### v1

@@ 4 Conclusion
In this work, we give a version of Monge’s and Kantorovich’s problems for lower probabilities. We then show that, for a particular class of cores of lower probabilities – the class of $\epsilon$-contaminations – our credal version of Monge’s problem coincides with the classical one, as does a restriction to our version of Kantorovich’s problem. We also give sufficient conditions for the existence of the lower optimal transport plan. Finally, we inspect when do our versions of the problems coincide, and find out that this need not hold in general. We also conjecture that this is a more general feature of credal sets, and not only of $\epsilon$-contaminations.
@@
With this work, we begin to explore the exciting venue of optimal transport between lower probabilities completely characterizing credal sets. In the future, we plan to further our study of optimal transport between $\epsilon$-contaminations by deriving a Brenier-type theorem (Brenier, 1991) and a Kantorovich-Rubinstein-type duality result (Galichon, 2016) (which could potentially have a significant impact in economics (Corrao et al., 2023; Kolotilin et al., 2023; Corrao and Dai, 2023; Corrao, 2023)). We also intend to explore the machine learning applications (especially concerning out of distribution detection) of our findings, and of distributionally robust optimization (Esfahani and Kuhn, 2018). We conjecture that the solutions to LPM (7) and RLPK (13) can bound the solutions to LPM and RLPK between two contaminated credal sets with different contamination parameters, i.e. $\mathcal{P}_{\epsilon}$ and $\mathcal{Q}_{\epsilon^{\prime}}$, $\epsilon^{\prime}\neq\epsilon$. Furthermore, we will extend our focus on other types of credal sets that are not necessarily completely characterized by lower probabilities, such as finitely generated credal sets (the convex hull of finitely many distributions), and to second-order distributions, that is, distributions over distributions.
Acknowledgements
We wish to express our gratitude to Yusuf Sale for insightful discussions, particularly on Lemma 2, Corollary 12.1, and Remark 1.

### v3

@@ 4 Conclusion
The conclusion that we can derive from this work is that Questions 1 and 2 in the Introduction have a positive answer. We can formulate a version of Monge’s and Kantorovich’s problems for lower probabilities. In addition, we can indeed find one class of credal sets completely characterized by their lower probability (the class of $\epsilon$-contaminations) for which the optimal transport map and plan coincide with the classical cases. We also inspected when our versions of the two problems coincide, and find out that this need not hold in general.
@@
With this work, we begin to explore the exciting venue of optimal transport between lower probabilities completely characterizing credal sets. In the future, we plan to further our study of optimal transport between $\epsilon$-contaminations by deriving a Brenier-type theorem (Brenier, 1991) and a Kantorovich-Rubinstein-type duality result (Galichon, 2016) (which could potentially have a significant impact in economics (Corrao et al., 2023; Kolotilin et al., 2023; Corrao and Dai, 2023; Corrao, 2023)). We also intend to explore the machine learning applications (especially concerning out-of-distribution detection) of our findings, and of distributionally robust optimization (Esfahani and Kuhn, 2018). Finally, we will extend our focus to other types of credal sets that are not necessarily completely characterized by lower probabilities, such as finitely generated credal sets (the convex hull of finitely many distributions), to signed measures, and to second-order distributions, that is, distributions over distributions.
Acknowledgments
We wish to express our gratitude to Yusuf Sale for insightful discussions, particularly on Lemma 2, Corollary 14.1, and Remark 1. We are also grateful to Krikamol Muandet and Siu Lun (Alan) Chau for suggesting the relationship between RLPK and the covariate shift condition in Machine Learning, and to four anonymous reviewers for their insightful comments.
Finally, as highlighted on the title page, we would like to dedicate the present work to the memory of Sayan Mukherjee. Without his guidance, we would never have had the strength – and perhaps the hubris – to start studying imprecise probability theory.
Appendix A On The Choice Of $\underline{P}$
In this section, we discuss our choice of the core $\mathcal{M}(\underline{P})$ in Lemma 6. We work with the (core of the incoherent, according to Walley (1991, Section 2.5)) lower probability $\underline{P}$ because it makes it easier to derive our desired results, e.g. the proof of Theorem 8.
Had we worked with $\underline{P}^{\prime}$ instead, we would have had (by Corollary 4.1)
$\int_{\mathcal{X}}c(x,T(x))\underline{P}^{\prime}(\text{d}x)=\int_{0}^{\infty}\underline{P}^{\prime}\left(\{x\in\mathcal{X}:c(x,T(x))\geq t\}\right)\text{d}t.$
It is not immediate to show that the latter is equal to $(1-\epsilon)\int_{\mathcal{X}}c(x,T(x)){P}(\text{d}x)$. To see this, notice that there might be some value $\bar{t}\in\mathbb{R}_{+}$ for which all $x\in\mathcal{X}$ are such that $c(x,T(x))\geq\bar{t}$. In that case, $\underline{P}^{\prime}\left(\{x\in\mathcal{X}:c(x,T(x))\geq\bar{t}\}\right)=\underline{P}^{\prime}(\mathcal{X})=1$, and so the “trick” that we used in (8) does not work anymore.
To achieve the desired result easily while working with $\underline{P}^{\prime}$, we would have had to require that the cost function $c$ is bounded. Indeed, suppose that the latter holds, and call $\underline{c}\coloneqq\inf_{x\in\mathcal{X}}c(x,T(x))$ and $\overline{c}\coloneqq\sup_{x\in\mathcal{X}}c(x,T(x))$. Then, we can use Troffaes and de Cooman (2014, Theorem C.3.(ii).(C.7)) to get
$\int_{\mathcal{X}}$ $c(x,T(x))\underline{P}^{\prime}(\text{d}x)$
$=\underline{P}^{\prime}(\mathcal{X})\cdot\underline{c}+\int_{\underline{c}}^{\overline{c}}\underline{P}^{\prime}\left(\{x\in\mathcal{X}:c(x,T(x))>t\}\right)\text{d}t$
$=\underline{c}+(1-\epsilon)\int_{\underline{c}}^{\overline{c}}{P}\left(\{x\in\mathcal{X}:c(x,T(x))>t\}\right)\text{d}t$
$=(1-\epsilon)\int_{\mathcal{X}}c(x,T(x)){P}(\text{d}x).$
As we can see, the desired result becomes either harder to prove (if we only ask for $c$ to be Borel measurable), or it needs an extra assumption (a bounded cost function $c$). A similar argument holds also for the Kantorovich’s results.
Since the main goal of the paper is to transport lower probabilities that completely characterize credal sets, and since by Lemma 6 we know that $\mathcal{M}(\underline{P})=\mathcal{M}(\underline{P}^{\prime})$, we opted for using the incoherent lower probability $\underline{P}$ instead of the coherent one $\underline{P}^{\prime}$.
We conclude with a remark. We acknowledge that working with the incoherent lower probability $\underline{P}$ makes it harder to use the techniques that we employ in this work, for models that are more complex than the $\epsilon$-contaminations that we study. How to overcome this shortcoming will be the object of future work.
Appendix B On The Difference Between Conditioning Methods
Let us illustrate the difference that the choice of conditioning rule makes when working with imprecise probabilities. Suppose that, instead of considering geometric, we choose generalized Bayes’ conditioning (Walley, 1991, Section 6.4). That is, for a generic credal set $\mathcal{P}$, for all $A\in\mathcal{F}$ and all $B\in\mathcal{G}$ such that $G(B)>0$,
$\underline{G}^{\text{GBC}}(A\mid B)\coloneqq\inf_{P\in\mathcal{P}}\left[\frac{G(A,B)}{G(B)}\right].$
By Walley (1991, Theorem 6.4.6), we have that
$\inf_{P\in\mathcal{P}}\left[\frac{G(A,B)}{G(B)}\right]=\frac{\underline{G}(A,B)}{\overline{G}(B)},$
so
$\underline{G}^{\text{GBC}}(A\mid B)$ $=\frac{\underline{G}(A,B)}{\overline{G}(B)}$
$\leq\frac{\underline{G}(A,B)}{\underline{G}(B)}\eqqcolon\underline{G}^{\text{geom}}(A\mid B),$
since $\overline{G}(B)\geq\underline{G}(B)$, for all $B\in\mathcal{G}$. This was also proven in Gong and Meng (2021, Lemma 4.3). This inequality still holds true even in a simple model like $\epsilon$-contaminated credal sets that we consider in the present work. To see this, notice that – in the same notation as Lemma 5 – by Wasserman and Kadane (1990, Example 3) in an $\epsilon$-contamination model $\mathcal{P}_{\epsilon}$ we have that $\overline{P}^{\prime}(A)=(1-\epsilon)P(A)+\epsilon$, for all $A\in\mathcal{F}\setminus\{\emptyset\}$, and $\overline{P}^{\prime}(\emptyset)=0$. Similarly to what we did in the main body of the paper, we can focus instead the incoherent upper probability $\overline{P}(A)=(1-\epsilon)P(A)+\epsilon$, for all $A\in\mathcal{F}$. Then, by Walley (1991, Section 6.6.2), we have that
$\underline{G}^{\text{GBC}}(A\mid B)=\frac{(1-\epsilon){G}(A,B)}{(1-\epsilon)Q(B)+\epsilon}$
and, similarly, that
$\underline{G}^{\text{GBC}}(B\mid A)=\frac{(1-\epsilon){G}(A,B)}{(1-\epsilon)P(A)+\epsilon}.$
In turn, this implies that the elements of $\Gamma_{R}^{\text{GBC}}(\underline{P},\underline{Q})$ are such that $\underline{G}(A,B)=\underline{G}^{\text{GBC}}(A\mid B)\overline{Q}(B)=\underline{G}^{\text{GBC}}(B\mid A)\overline{P}(A)$, for all $A\in\mathcal{F}$ and all $B\in\mathcal{G}$ (having positive upper probability). They are different than the elements of $\Gamma_{R}^{\text{geom}}(\underline{P},\underline{Q})$ that we introduced in Definition 10.
Appendix C Proof of Lemma 2
[…]
Notice that $\mathcal{M}^{\text{bc}}(T_{\#}\underline{P})$ is nonempty because it contains the pushforward measures of the elements of $\mathcal{M}^{\text{fa}}(\underline{P})$ in (1) through map $T$. In formulas, $P\in\mathcal{M}^{\text{fa}}(\underline{P})\implies T_{\#}P\in\mathcal{M}^{\text{bc}}(T_{\#}\underline{P})$. To see that this is the case, notice that a finitely additive probability is a special case of a bounded charge, and that $\mathcal{M}^{\text{fa}}(\underline{P})$ contains all the finitely additive probabilities on $\mathcal{X}$ that set-wise dominate $\underline{P}$. More formally, pick any $\tilde{B}\in\mathcal{G}$ and any $P\in\mathcal{M}^{\text{fa}}(\underline{P})$. Let $\tilde{A}=T^{-1}(\tilde{B})$. Then, by (1), $T_{\#}P(\tilde{B})=P(T^{-1}(\tilde{B}))=P(\tilde{A})\geq\underline{P}(\tilde{A})=\underline{P}(T^{-1}(\tilde{B}))=T_{\#}\underline{P}(\tilde{B})$. But then $T_{\#}P\in\mathcal{M}^{\text{bc}}(T_{\#}\underline{P})$, which shows that $\mathcal{M}^{\text{bc}}(T_{\#}\underline{P})\neq\emptyset$. In turn, by Marinacci and Montrucchio (2004, Proposition 3), we have that $\mathcal{M}^{\text{bc}}(T_{\#}\underline{P})$ is weak-compact.
Now, notice that $\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})\coloneqq\{Q\in\Delta^{\text{fa}}_{\mathcal{Y}}:Q(B)\geq T_{\#}\underline{P}(B)\text{, }\forall B\in\mathcal{G}\}$ is a proper subset of $\mathcal{M}^{\text{bc}}(T_{\#}\underline{P})$, since $\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})$ only considers the finitely additive probabilities (and not all the bounded charges) that set-wise dominate $T_{\#}\underline{P}$.
Let now $(Q_{\alpha})_{\alpha\in\mathcal{I}}$ be a net in $\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})$ that weak-converges to $Q\in\Delta^{\text{fa}}_{\mathcal{Y}}$. Here $\mathcal{I}$ is a generic index set. This means that $Q_{\alpha}(B)\rightarrow Q(B)$, for all $B\in\mathcal{G}$. That is, pick any $B\in\mathcal{G}$; then,
Equation (20) implies that, for all $Q_{\alpha}\succeq Q_{\tilde{\alpha}_{\epsilon}}$, we have that $Q(B)>Q_{\alpha}(B)-\epsilon\geq T_{\#}\underline{P}(B)-\epsilon$. Letting $\epsilon\rightarrow 0$, this implies that $Q(B)\geq T_{\#}\underline{P}(B)$. But $B$ was chosen arbitrarily in $\mathcal{G}$, and so $Q\in\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})$. Hence, $\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})$ is weak sequentially closed, and therefore it is weak closed. Being a weak closed subset of a weak compact space, we can conclude that $\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})$ is weak compact itself.
By Cerreia-Vioglio et al. (2015, Section 2.1.(viii)), we know that a (set) function $\nu:\mathcal{G}\rightarrow[0,1]$ is a lower probability if and only if there exists a weak-compact set $\mathcal{M}\subseteq\Delta^{\text{fa}}_{\mathcal{Y}}$ such that $\nu(B)=\min_{Q\in\mathcal{M}}Q(B)$, for all $B\in\mathcal{G}$. Letting $\mathcal{M}\equiv\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})$ and $\nu\equiv T_{\#}\underline{P}$, we obtain that $T_{\#}\underline{P}(B)\leq\min_{Q\in\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})}Q(B)$, for all $B\in\mathcal{G}$. Suppose now for the sake of contradiction that there is $\tilde{B}\in\mathcal{G}$ such that $T_{\#}\underline{P}(\tilde{B})<\min_{Q\in\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})}Q(\tilde{B})$. This would imply that $\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})\supsetneq\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})$, a contradiction. Hence, we can conclude that $T_{\#}\underline{P}(B)=\min_{Q\in\mathcal{M}^{\text{fa}}(T_{\#}\underline{P})}Q(B)$, for all $B\in\mathcal{G}$, thus completing the proof. ∎